### v4

# CrackSCF: Lightweight Cascaded Fusion Network for Robust and Efficient Structural Crack Segmentation

Abstract
Accurately segmenting structural cracks at the pixel level remains a major hurdle, as existing methods fail to integrate local textures with pixel dependencies, often leading to fragmented and incomplete predictions. Moreover, their high parameter counts and substantial computational demands hinder practical deployment on resource-constrained edge devices. To address these challenges, we propose CrackSCF, a Lightweight Cascaded Fusion Crack Segmentation Network designed to achieve robust crack segmentation with exceptional computational efficiency. We design a lightweight convolutional block (LRDS) to replace all standard convolutions. This approach efficiently captures local patterns while operating with a minimal computational footprint. For a holistic perception of crack structures, a lightweight Long-range Dependency Extractor (LDE) captures global dependencies. These are then intelligently unified with local patterns by our Staircase Cascaded Fusion Module (SCFM), ensuring the final segmentation maps are both seamless in continuity and rich in fine-grained detail. To comprehensively evaluate our method, we created the challenging TUT benchmark dataset and evaluated it alongside five other public datasets. The experimental results show that the CrackSCF method consistently outperforms the existing methods, and it demonstrates greater robustness in dealing with complex background noise. On the TUT dataset, CrackSCF achieved 0.8382 on F1 score and 0.8473 on mIoU, and it only required 4.79M parameters. The code and dataset is available at https://github.com/Karl1109/CrackSCF.
{IEEEkeywords}
Crack segmentation, lightweight networks, cascaded fusion, local patterns, long-range dependencies
1 Introduction
[…]
Cracks are common defects in industrial structures such as pavements and buildings, which can lead to serious safety issues. Thus, automated detection on edge control devices is crucial for ensuring production safety [1]. Early automated detection methods used traditional image processing techniques for crack extraction [2]. While these methods are easy to implement, their performance in complex scenarios still needs improvement.
Figure 1: (a) shows the performance of our method compared to other methods. (b) shows the different scenarios included in TUT dataset.
In recent years, benefiting from the strong local inductive capability of Convolutional Neural Networks (CNN), CNN-based networks such as Crack500 [3] and SC-OSDA [4] have shown efficient feature extraction capabilities in structural crack segmentation tasks. The spatial invariance of CNNs allows them to detect cracks anywhere in an image by learning similar features at different locations. In labeled images, crack labels are long, narrow structures of white pixels, while the background is large black areas. This challenges the model to distinguish between thin cracks and the background. Nevertheless, networks based on CNNs face challenges in capturing long-range dependencies between crack pixels, leading to segmentation discontinuities.
The success of Vision Transformer (ViT) [5] has demonstrated the advantages of Transformer, with their strong sequence processing capabilities, in handling visual tasks [6]. Unlike CNNs, transformers can capture dependencies between any pixels when processing images, making them more advantageous in handling complex topologies. To harness the advantages of both CNNs and transformers effectively, researchers have proposed networks such as DTrCNet [7] and Crackmer [7], which integrate local and long-range information through U-shaped or residual connections.
Despite the aforementioned methods combining CNN and Transformer have improved performance in pixel-level crack segmentation tasks, they still have some limitations, as follows:
  1. 1.
    The different fusion strategies significantly affect the segmentation performance of CNN-Transformer networks. For instance, UCTNet [8] and CTCS [9] use channel concatenation to fuse local patterns and long-range dependences information. However, concatenation merely stacks the feature information together without deeper feature interaction, failing to understand their correlation and complementarity, resulting in poor segmentation performance on fine cracks and susceptibility to background noise. Additionally, DTrCNet [7] and Crackmer [10] use channel attention to fuse information, which can enhance focus on certain features but neglects spatial information and pixel details in the feature maps, impacting the model’s overall performance.
  2. 2.
    Most networks, in pursuit of performance, overlook the parameter and computational costs brought by numerous convolution operations, this makes them challenging to deploy on resource-constrained control devices. As shown in Fig. 1(a), networks like SFIAN [11], Crackmer [10], and CTCS[9] have high computational parameter counts, indicating that they are difficult to implement on resource-limited devices like smartphones and drones, significantly diminishing their practicality.
To address these issues, we propose a staircase cascaded fusion network that generates high-quality segmentation maps with low computational resources while effectively suppressing background noise. Specifically, to reduce the number of parameters and computational burden, we introduce a lightweight convolution block called LRDS. This module utilizes the principle of low-rank approximation to effectively reduce the computational resources required by convolutional layers, and captures spatial structures efficiently in the low-rank space through depthwise convolution, maintaining good feature perception while requiring very low computational resources. We also propose a lightweight Long-range Dependency Extractor (LDE), which uses a lightweight deformable attention mechanism to effectively capture irregular long-range dependencies between crack pixels with very low computational demands. Additionally, to better leverage the complementarity and correlation of feature maps while ensuring that critical detail information is not missed, we introduce the Staircase Cascaded Fusion Module (SCFM). This module generates feature maps layer by layer and effectively fuses them using a pixel attention mechanism that can finely capture local and long-range information and a channel concatenation mechanism, comprehensively combining local texture information and pixel dependencies.
To comprehensively evaluate the structural crack segmentation performance of the model in various complex environments, we also collected a dataset named TUT, which contains 1408 RGB images. Existing public datasets like DeepCrack [12] and Crack500 [3] include only 1-2 types of scenarios and have relatively simple crack shapes, which are insufficient for thoroughly testing the network’s performance. As shown in Fig. 1(b), the TUT dataset includes a broader range of real structural crack image scenarios, covering plastic runways, bricks, tiles, cement, bitumen, generator blades, metal materials, and underground pipelines. With such a diverse set of image scenarios, this dataset effectively evaluates the network’s generalization and robustness. Furthermore, the images feature occlusions, highly complex backgrounds, and uneven lighting, which further enhances the datasets diversity.
In conclusion, our main contributions are as follows:
  1. 1.
    We proposed a novel CrackSCF network for structural crack segmentation. By introducing a lightweight convolution block, a lightweight long-range dependency extractor, and a staircase cascaded fusion module, this network effectively distinguishes between background and crack pixels with low computational resources, resulting in high-quality segmentation maps.
  2. 2.
    We collected the TUT dataset, which includes structural cracks in eight different scenarios, providing a more comprehensive evaluation of the network’s performance.
  3. 3.
    We carried out comprehensive experiments on five public datasets and the TUT benchmark dataset, demonstrating that our proposed CrackSCF network outperforms existing SOTA methods.
The organization of the subsequent sections of this paper is as follows: In Section 2, existing crack detection models and lightweight networks are reviewed. Section 3 details the components of the proposed network. Section 4 delves into the experimental details and results. Finally, Section 5 provides a summary of the entire paper and highlights future research directions.
2 Related Work
Figure 2: General Architecture Diagram of CrackSCF Network. Crack images are first input into the MFE, generating four feature maps. After enhancement, five feature maps with unified channel numbers are obtained. The LDE processes these maps to acquire pixel sequences rich in long-range correlations. The SCFM then processes these sequences and the four layers of feature maps from the MFE. Through four stages, resulting in the segmentation output.
2.1 Crack Segmentation Networks
Recently, a growing number of researchers have utilized CNNs to develop crack segmentation networks, achieving advanced performance. Zhu et al. [13] proposed the BSNBM network, which integrates bidirectional self-rectifying modules with Bayesian inference to enhance crack detection robustness and accuracy under uncertainty. Cheng et al. [11] introduced the Selective Feature Fusion and Irregular-Aware Network (SFIAN) to effectively model irregular crack objects, resulting in clearer backgrounds on various crack images. Lyu et al. [14] proposed a CNN-based crack quantification framework that utilizes voxel-based reconstruction and Bayesian data fusion to achieve accurate crack width estimation.
Benefit from the attention mechanism, transformers are particularly well-suited for handling targets with long-range dependencies and complex global structures, especially for cracks, which often have extensive spans and intricate shapes. Therefore, many researchers are amalgamating the outputs of CNNs and transformers for crack segmentation. Qi et al. [15] developed an end-to-end model based on ViT and level set theory for segmenting defects in bridge pavement, achieving precise segmentation by amalgamating the outputs of two parallel decoders. Feng et al. [16] developed a shale crack segmentation network that combines CNN and Swin Transformer, utilizing Fast Fourier Transform (FFT) to merge feature maps from both, effectively suppressing high-frequency noise in the imaging process.
Although CNN and CNN-Transformer methods have shown promise in crack segmentation, CNNs struggle to capture global context due to their local receptive fields, leading to discontinuous predictions and vulnerability to background noise. CNN-Transformer hybrids improve global perception but often adopt single-branch fusion, limiting their ability to integrate local and long-range cues. Moreover, both lack selective interaction and noise suppression, and their high computational cost hinders deployment on edge devices.
2.2 Lightweight Networks
Deploying deep learning models on edge devices like smartphones and drones is trending. However, these devices have limited computing resources, making it challenging to use models with high parameter counts. Thus, compact model design for lightweight modifications in size and speed is crucial. For example, Howard et al. [17] proposed MobileNet, an efficient model using depthwise separable convolutions to reduce computational load and parameter count, suitable for resource-constrained devices. Ma et al. [18] introduced ShuffleNet, providing practical design principles for efficient network architecture. Additionally, low-rank approximation compression, effective in fields like numerical linear algebra and image processing, is now used to compress deep learning network architectures. Swaminathan et al. [19] introduced SLR, a sparse low-rank method that improves compression rates with sparse SVD matrices.
In recent years, there has been a growing trend among researchers to integrate model lightweighting techniques with crack segmentation models, aiming to facilitate real-time crack detection on edge computing devices. For example, Zhu et al. [20] designed RHACrackNet, a lightweight crack detection network integrating hybrid attention and residual blocks, enabling real-time detection on edge devices. Huang et al. [21] proposed a lightweight fusion attention module for road crack detection, ensuring high-precision crack extraction with rapid detection. Although lightweight models reduce computational cost, their aggressive parameter reduction limits model expressiveness, impairing the capture of intricate patterns and leading to suboptimal segmentation. This trade-off between performance and efficiency undermines their reliability in real-world applications.
3 Methodology
3.1 Network Architecture
As shown in Fig. 2, the CrackSCF network has three main components: a lightweight Multi-scale Feature Extractor (MFE), a lightweight Long-range Dependency Extractor (LDE), and a Staircase Cascaded Fusion Module (SCFM). The overall architecture of the network is shaped like a staircase, with feature maps of different scales progressively fused and upsampled. In this process, the advantages of CNNs and Transformers in feature extraction complement each other.
Specifically, for a given dataset, the images are first input into the MFE, which is divided into four stages. In each stage, the extracted feature maps undergo a Feature Enhancement stage. Then, the LDE processes these enhanced feature maps to obtain pixel sequences rich in long-range dependencies. Subsequently, the SCFM receives these sequences along with the four layers of feature maps extracted by the MFE. After processing through four stages, the resolution is doubled at each stage while the number of channels is halved, ultimately resulting in the segmentation result.
Importantly, as shown in Fig. 3, using the original convolution operations results in high FLOPs and Params, especially in the SCFM module, reaching 73.36G and 9.71M, respectively. Replacing all convolutions with LRDS blocks significantly reduced FLOPs and Params, demonstrating the effectiveness of LRDS in model lightweighting.
Figure 3: Comparison of each module when using our LRDS or original convolution separately. (a) shows the comparison of FLOPs, (b) shows the comparison of parameters.
3.2 Lightweight Multi-scale Feature Extractor
As shown in Fig. 2, our designed LRDS block adopts a bottleneck-like structure. It comprises a Reduce layer, a depthwise convolution layer, a pointwise convolution layer, and an Expand layer. The input feature map first passes through the Reduce layer, reducing the input channels to $f_{mid}$. It then goes through the depthwise convolution layer and the pointwise convolution layer, keeping the number of channels unchanged. Finally, the Expand layer increases the number of channels to $f_{out}$, producing the final output feature map.
Low-rank approximation decomposes a high-dimensional matrix into low-dimensional matrices, reducing computation complexity. In CNNs, it is applied to convolutional kernel responses, assuming these lie in a low-rank subspace.
Specifically, in the convolutional layer, we assume that its filter has a spatial size of $k$ and a number of input channels of $c$. In computing the convolutional response, the filter is applied to a sub-tensor of the input tensor of shape $k\times k\times c$. Suppose this sub-tensor is $t\in\mathbb{R}^{k^{2}c+1}$, where we append a bias term at the end. The response at a position in the output, $y\in\mathbb{R}^{e}$, can be expressed as follows:
$y=Wt$ (1)
where $e$ represents the number of filters and $W$ is a $e\times(k^{2}c+1)$ matrix. Assuming the vector $y$ lies in a low-rank subspace, we can write it as $y=U(y-y_{1})+y_{1}$, where $U$ is a $e\times e$ matrix of rank $e_{0}$, with $e_{0}<e$. $y_{1}$ is the mean vector of the response. Therefore, we can write the response as
$y=UWt+b$ (2)
where $b=y_{1}-Uy_{1}$ is the new bias. The rank of $U$ is $e_{0}$, and it can be decomposed into two $e\times e_{0}$ matrices $J$ and $K$ such that $U=JK^{\top}$. We use $W_{1}=K^{\top}W$ to denote a $e_{0}\times(k^{2}c+1)$ matrix, which corresponds to a set of $e_{0}$ filters. Thus, we can rewrite the response as follows:
$y=JW_{1}t+b$ (3)
The computational complexity of using Equation (3) is $O(e_{0}k^{2}c)+O(ee_{0})$, whereas approximating $y$ in the low-rank subspace using Equation (1) has a computational complexity of $O(ek^{2}c)$. Given that $O(ee_{0})\ll O(e_{0}k^{2}c)$, the complexity of using Equation (3) is $\frac{e_{0}}{e}$ times that of using Equation (1).
In order to efficiently capture the spatial information of feature maps in low-rank space for finer perception of detailed regions, we process the features in LRDS by embedding deep convolution and pointwise convolution in low-rank space. We constructed MFE by replacing all convolution operations with LRDS convolution blocks.
3.3 Lightweight Long-range Dependency Extractor
Figure 4: Analysis of Computational Resources Required for Linear Operations. (a) shows the percentage of FLOPs and Params in the LDE module that use Original linear operations, with the inner circle showing the percentage that use FLOPs and the outer circle showing the percentage of Params. (b) Shows the comparison of FLOPs and Params for LDE before and after using LR Linear.
When extracting long-range dependencies of crack pixels, the deformable attention mechanism adapts well to irregular structures. Specifically, given an input $x\in\mathbb{R}^{C\times H\times W}$, let $q$ denote a query element with query feature $z_{q}$ and reference point coordinates $(p_{qx},p_{qy})$, the deformable attention is calculated as:
$\centering\begin{split}&MSDeformAttn(z_{q},\hat{p}_{q},\{x_{c}\}_{c=1}^{C})=\\ &\sum_{h=1}^{H}W_{h}\Bigg[\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}\cdot W^{\prime}_{h}x_{c}(\phi_{c}(\hat{p}_{q})+\Delta p_{hcqt})\Bigg]\end{split}\@add@centering$ (4)
where $h$ represents the attention head, $c$ represents the input feature layer, and $t$ represents the sampling key point. $\Delta p_{hcqt}$ and $A_{hcqt}$ represent the offset and attention weight of the $t$-th sampling point in the $c$-th feature map and $h$-th attention head, respectively. The normalization is such that $\sum_{c=1}^{C}\sum_{t=1}^{T}A_{hcqt}=1$. Here, normalized coordinates $\hat{p}\in[0,1]$ are used to represent the position of a reference point in the normalized feature map $\phi_{c}(\hat{p}_{q})$ rescales the normalized coordinates onto the $c$-th layer feature map.
As shown in Fig. 4(a), Linear operations account for 90% of the FLOPs and 67% of the Params in the LDE module. In order to reduce the computational resources required by the LDE, we propose a linear transformation method (LR Linear) based on low-rank approximation. In traditional linear layers, the weight matrix connecting the input and output is typically a high-rank matrix, representing a complete connection between input and output channels. By applying low-rank approximation, we decompose the weight matrix into two smaller matrices, reducing the number of parameters.
Specifically, the weight matrix $E\in\mathbb{R}^{r_{\text{out}}\times r_{\text{in}}}$, where $r_{\text{out}}$ is the output feature dimension and $r_{\text{in}}$ is the input feature dimension, is approximated by two smaller matrices $A\in\mathbb{R}^{r_{\text{out}}\times r}$ and $B\in\mathbb{R}^{r_{\text{in}}\times r}$, where $r$ is a small rank, much smaller than $r_{\text{out}}$ and $r_{\text{in}}$. This decomposition breaks the original full connection into two smaller linear transformations:
$E\approx E^{\prime}=AB^{T}$ (5)
we perform this transformation through two linear layers. First, we apply a linear transformation $B^{T}$ to the input $x\in\mathbb{R}^{r_{\text{in}}}$, resulting in an intermediate representation $z\in\mathbb{R}^{r}$:
$z=B^{T}x$ (6)
Then, we apply another linear transformation $A$ to $z$, resulting in the output $y\in\mathbb{R}^{r_{\text{out}}}$:
$y=Az$ (7)
As $r\ll r_{\text{out}},r_{\text{in}}$, the low-rank approximation reduces the original matrix to two smaller matrices, significantly reducing the computational and storage requirements. This leads to a reduction in both the computational complexity and the number of parameters, thus improving training and inference efficiency. As shown in Fig. 4(b), replacing all Linear operations in the LDE with LR Linear reduced the parameters and computational load from 7.18G and 1.05M to 0.14G and 0.02M.
The structure of LDE is shown in Fig. 2. LDE facilitates long-range dependency modeling through the interaction of multi-scale features between the encoder and decoder. Initially, the encoder inputs a multi-scale feature set $\{X_{c}\}_{c=1}^{C}$ (where $C$ represents the number of feature levels). After undergoing layer flattening and positional encoding enhancement, the encoder extracts the feature sequence $S\in\mathbb{R}^{B\times L\times d}$ (where $L$ denotes the sequence length). The encoder then combines LR Linear with lightweight deformable attention mechanisms to capture long-range dependencies:
$Z=MSDeformAttn(S,P,S)$ (8)
where $P\in\mathbb{R}^{B\times L\times 2}$ represent the normalized reference coordinates. Subsequently, $Z$ is processed through a feedforward network (FFN) that combines LR Linear to refine the features:
M=LR_Linear(GELU(LR_Linear(Z)))M=LR\_Linear(GELU(LR\_Linear(Z))) (9)
where the output memory matrix $M\in\mathbb{R}^{B\times L\times d}$ is used as the input to the decoder, encoding multi-scale local dependencies.
The decoder processes the input with a learnable query vector $Q\in\mathbb{R}^{N\times d}$ and the initial reference point $H=\text{Sigmoid}(TQ)$, where $T\in\mathbb{R}^{d\times 2}$ is the projection matrix. First, the reference points $H\in\mathbb{R}^{N\times 2}$ are mapped to the encoded memory $M$ to aggregate features:
$Y_{0}=MSDeformAttn(Q,H,M)$ (10)
Then, it passes through the FFN to obtain the crack pixel sequence enriched with irregular long-range dependencies:
$Y=LR\_Linear(GELU(LR\_Linear(Y_{0})))$ (11)
where Y is the output sequence of the LDE enriched with irregular long-range dependencies between pixels.
3.4 Staircase Cascaded Fusion Module
The design of our SCFM structure is shown in Fig. 2. As a segmentation head, to enhance useful semantic features in the MFE and LDE branches and ensure effective semantic information is not buried, we proposed the lightweight Pixel Attention Guided Fusion (PAF) module. The structure of PAF is shown in Fig. 2. We denote the pixels in the feature maps of the two branches as $\vec{v_{c}}$ and $\vec{v_{d}}$, respectively. The output can be expressed as:
$Output=\sigma\vec{v_{c}}+(1-\sigma)\vec{v_{d}}$ (12)
where $\sigma$ indicates the probability that the two pixels belong to the same category. If $\sigma$ is high, the model will trust $\vec{v_{c}}$ more, indicating that its semantic features are more accurate, and vice versa.
To reduce computational resource usage, all convolution operations in this module utilize LRDS blocks. As seen in Fig. 3, using original convolution operations in the SCFM results in FLOPs and Params reaching 73.36G and 9.71M, respectively. Thus, lightweighting these operations is crucial. Replacing all convolution operations with LRDS blocks reduces FLOPs and Params by 92.87% and 97.22%, respectively, demonstrating the effectiveness of LRDS in network lightweighting.
SCFM is divided into four stages, each with two fusion branches: the concat branch and the PAF branch. The concat operation retains complete information and diverse feature representations through direct concatenation, while the PAF module selectively fuses features through a pixel-guided attention mechanism, avoiding issues like buried feature information, information conflict, and redundancy. The processes of SCFM can be given by the following equation:
$\begin{split}&F_{i}^{O}=C_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\\ &+\mu\left(P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\right)\odot P_{i}\left(F_{i}^{MFE},F_{i-1}^{O}\right)\end{split}$ (13)
Where $F_{i}^{\text{O}}$ represents the output features of the $i$-th SCFM layer ($i$ ranges from 0 to 3), $F_{i}^{\text{MFE}}$ represents the $i$-th layer MFE features, $C_{\text{i}}$ is an operation chain: Concat followed by LRDS, ReLU, and GN, $P_{i}$ is a PAF operation, $\mu(\cdot)$ is the sigmoid function generating spatial attention weights, and $\odot$ represents the Hadamard product. Specifically, when $i=0$, $F_{i-1}^{\text{O}}$ is defined as the output of LDE, i.e., $F_{-1}^{\text{O}}=\text{LDE}_{\text{output}}$, ensuring an initial output feature is provided at the beginning of the computation.
By perceiving the feature maps from the two branches, both local patterns of crack pixels and long-range dependencies can be comprehensively addressed. To fully utilize the outputs of MFE and LDE and improve segmentation performance, SCFM utilized the property that PAF can process multiple feature maps simultaneously by merging the outputs of all the preceding phases and the associated MFE feature maps from the second phase onwards. The processing enables SCFM to more comprehensively dissect and process the local patterns and long-range dependencies in the different stages of the feature map to achieve a finer and more continuous segmentation map. Furthermore, the fine-grained features extracted with PAF can be perceived more finely through the dynamic weighting mechanism to achieve a clearer segmentation map.
3.5 Loss Function
The discrepancy between predictions and actual values is quantified by the loss function; a smaller value for this loss suggests superior model performance. Previously, Binary Cross-Entropy (BCE) was often used, but it may not effectively distinguish the minority crack pixels in the image. Combining BCE with Dice loss addresses this issue. BCE improves the probability distribution for each pixel, while Dice loss maximizes the overlap between predictions and true labels, enhancing robustness against imbalanced data [22] .
Our loss $L$ can be calculated using Equation (14), which combines the above two loss functions for joint optimization.
$L=0.75\cdot L_{B}+0.25\cdot L_{D}$ (14)
where $\mathit{L_{B}}$ represents the BCE loss function and $\mathit{L_{D}}$ represents the Dice loss function.
4 Experiments
In this section, we analyze the differences between our TUT dataset and other public datasets, describe the experimental setup, and compare methods and metrics. We then analyze results and model complexity, perform ablation experiments.
4.1 Dataset Description
Table 1: Comparison of Different Crack Datasets
No. Datasets Year Num(train/val/test) Resolution Percentage of cracked pixels Scenarios
1 CFD [23] 2016 118(76/11/31) 480 width and 320 height 1.60% bitumen
2 GAPS509 [24] 2017 509(307/100/102) 540 width and 640 height 1.19% bitumen
3 DeepCrack [12] 2019 537(300/100/247) 544 width and 384 height 3.54% bitumen, cement
4 Crack500 [3] 2019 3368(1896/348/1124) 640 width and 360 height 6.01% bitumen, cement
5 CrackMap [25] 2023 120(72/12/36) 256 width and 256 height 4.10% bitumen
6 TUT (Ours) 2024 1408(987/139/282) 640 width and 640 height 3.16% bitumen, cement, bricks, runways, metal, blades, pipelines, tiles
The previous works has used dataset scenarios that are relatively simple, such as cracks in cement or bitumen pavements only. These datasets have clear backgrounds and minimal interference, which cannot comprehensively evaluate the performance of models. To tackle this problem, we created the TUT dataset to enable better segmentation across different scenarios. Detailed information and comparisons with other datasets are list in Table 1.
The TUT dataset comprises 1270 crack images, along with a small number of 138 images sourced from the Internet. TUT dataset contains 900 images of bitumen and cement with irregular crack morphology but flat backgrounds, 142 images of bricks and tiles with regular crack morphology, 228 images of plastic runways with extremely complex cracks, and a total of 138 images of metals, generator blades, and underground pipelines with complex lighting and background disturbances. To ensure reliability in model training and testing, all images were manually annotated by several researchers to generate binary labels, and the best-annotated images were selected as the final labeled images. Unlike other datasets with clear backgrounds, the TUT dataset images have complex backgrounds and cracks of intricate shapes. This evaluates the model’s capability to segment cracks in challenging conditions.
Figure 5: Illustration of the various scenarios of the TUT dataset with the original images and Ground Truth.
As list in Table 1, our TUT dataset contains eight image scenarios, far more than the 1-2 scenarios of other public datasets, this allows for a more comprehensive evaluation of the model’s performance in detecting cracks under various scenarios. Fig. 5 shows crack images in different scenarios and their binary labeled images. The crack pixel ratio in the TUT dataset is 3.16%, which is moderate. This ensures the model learns useful information without being hindered by too few crack pixels and prevents the model from overly relying on the crack pixel ratio when there are too many crack pixels.
As seen in Fig. 5, bitumen, cement, bricks, and plastic runways have extremely complex cracks due to their unique material properties. Particularly in plastic runways, both coarse and fine cracks are present, testing the model’s generalization to different types of cracks. In images of tiles, metal materials, wind turbine blades, and underground pipelines, the crack shapes are relatively simple, but the backgrounds are very complex with various external interferences such as low light, irrelevant areas, and surface characteristics. This can easily result in false detections, thereby testing the model’s ability to accurately extract crack areas from complex environments.
4.2 Implementation Details
1) Experimental Settings: The CrackSCF network is built on the PyTorch v2.1.1 deep learning framework and trained on an Ubuntu 20.04 server, featuring two Intel(R) Xeon(R) Platinum 8336C CPUs and eight Nvidia GeForce 4090 GPUs, each with 24GB VRAM. For training, the AdamW optimizer is utilized with an initial learning rate set at 0.0005, weight decay of 0.0001, batch size of one, and a total of 60 training epochs. The learning rate is reduced to one-tenth after 30 epochs as part of a decay strategy. The best-performing model parameters on the validation set are selected for testing. We compared the CrackSCF network with other SOTA methods on six datasets: CFD [23], GAPS509 [24], DeepCrack [12], Crack500 [3], CrackMap [25], and ours TUT dataset.
Table 2: Comparison of results between CrackSCF and other methods on the public dataset and the TUT dataset. Best results are bolded and second best results are underlined
Methods CFD GAPS509 Crack500
ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU ODS OIS P R F1 mIoU
RIND 0.6858 0.7021 0.6456 0.7534 0.6954 0.7592 0.5505 0.6023 0.4959 0.6494 0.5623 0.6909 0.6469 0.6483 0.6998 0.7245 0.7119 0.7381
UCTNet 0.6587 0.7006 0.6701 0.7205 0.6944 0.7479 0.5835 0.6188 0.5709 0.6752 0.6187 0.7107 0.7147 0.7265 0.6887 0.7843 0.7302 0.7674
SFIAN 0.7040 0.7159 0.6657 0.7512 0.7058 0.7684 0.6087 0.6320 0.6061 0.6451 0.6250 0.7198 0.6977 0.7348 0.6983 0.7742 0.7343 0.7604
CTCS 0.7122 0.7179 0.6896 0.7438 0.7157 0.7738 0.6029 0.6075 0.5693 0.6804 0.6199 0.7165 0.6941 0.7059 0.6940 0.7748 0.7322 0.7591
DTrCNet 0.7079 0.7123 0.6774 0.7486 0.7112 0.7710 0.6211 0.6276 0.6277 0.6592 0.6431 0.7272 0.7012 0.7241 0.6527 0.8280 0.7357 0.7627
Crackmer 0.6751 0.6754 0.6370 0.7237 0.6776 0.7503 0.4593 0.4642 0.6144 0.4663 0.5302 0.6581 0.6933 0.7097 0.6985 0.7572 0.7267 0.7591
SimCrack 0.7152 0.7204 0.6883 0.7557 0.7204 0.7756 0.6408 0.6471 0.6171 0.6997 0.6558 0.7368 0.7127 0.7308 0.7093 0.7984 0.7516 0.7715
Ours 0.7335 0.7376 0.6967 0.7848 0.7383 0.7864 0.6492 0.6552 0.6459 0.6812 0.6631 0.7411 0.7226 0.7353 0.7213 0.7922 0.7551 0.7764
CrackMap DeepCrack TUT (Ours)
RIND 0.6745 0.6943 0.6023 0.7586 0.6699 0.7425 0.8087 0.8267 0.7896 0.8920 0.8377 0.8391 0.7531 0.7891 0.7872 0.7665 0.7767 0.8051
UCTNet 0.6104 0.6195 0.5653 0.6748 0.6152 0.7079 0.8357 0.8579 0.8217 0.8857 0.8525 0.8564 0.7569 0.7670 0.8201 0.7667 0.7925 0.8110
SFIAN 0.7200 0.7465 0.6715 0.7668 0.7160 0.7748 0.8616 0.8928 0.8549 0.8692 0.8620 0.8776 0.7290 0.7513 0.7715 0.7367 0.7537 0.7896
CTCS 0.7289 0.7373 0.6911 0.7669 0.7270 0.7785 0.8819 0.8904 0.9011 0.8895 0.8952 0.8925 0.7940 0.7996 0.8202 0.8195 0.8199 0.8301
DTrCNet 0.7328 0.7413 0.6912 0.7681 0.7276 0.7812 0.8473 0.8512 0.8905 0.8251 0.8566 0.8661 0.7987 0.8073 0.7972 0.8441 0.8202 0.8331
Crackmer 0.7395 0.7437 0.7229 0.7467 0.7346 0.7860 0.8712 0.8785 0.8946 0.8783 0.8864 0.8844 0.7429 0.7640 0.7501 0.7656 0.7578 0.7966
SimCrack 0.7559 0.7625 0.7380 0.7672 0.7523 0.7963 0.8570 0.8722 0.8984 0.8549 0.8761 0.8744 0.7984 0.8090 0.8051 0.8371 0.8208 0.8334
Ours 0.7595 0.7651 0.7390 0.7708 0.7546 0.7987 0.8914 0.8963 0.9147 0.9013 0.9079 0.9001 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473
2) Comparison Methods: To evaluate the model performance, we compared the CrackSCF with seven classical and latest SOTA methods on different datasets. These methods include RIND [26], UCTNet [8], SFIAN [11], CTCS [9], DTrCNet [7], Crackmer [10] and Simcrack [27].
3) Evaluation Metric: To quantitatively evaluate the performance of our proposed crack segmentation model, we used six main metrics [11]: Precision (P), Recall (R), F1 Score ($F1=\frac{2PR}{P+R}$), mean Intersection over Union (mIoU), Optimal Dataset Scale (ODS), and Optimal Image Scale (OIS).
4.3 Comparison with the SOTA methods
1) Results on the our TUT dataset: Fig. 6 shows the visual results of all methods on the TUT dataset. We manually highlighted easily missed crack pixel areas with red boxes and misdetected non-crack pixel areas with yellow boxes. As list in Table 2, our method achieved excellent results on the TUT dataset, leading in all metrics. Specifically, it outperformed the second-best methods on ODS, OIS, P, R, F1, and mIoU by 2.73%, 2.18%, 0.98%, 0.51%, 2.12%, and 1.67%, respectively. SimCrack [27] achieved the second-best results on ODS, OIS, F1, and mIoU, but as seen in Fig. 6, its segmentation performance on some key details has room for improvement compared to our method.
Furthermore, most methods fail to suppress irrelevant background noise. In images of metal materials and turbine blades, our method effectively removes noise and accurately segments crack regions. In pipeline scenarios, our method can filter out timestamp watermarks, thanks to SCFM, which enables the model to focus on irregular, slender crack structures rather than irrelevant watermarks. Our method demonstrates excellent robustness and versatility, resulting in high-quality crack segmentation maps and adapting well to different scenarios.
2) Results on other public datasets: The CFD, CrackMap and GAPS509 datasets are characterized by slender and fine-grained crack structures, testing the model’s sensitivity to subtle and narrow crack pixels. As listed in Table 2, our proposed CrackSCF achieved the best performance on the CFD and CrackMap datasets. On both datasets, CrackSCF outperforms Simcrack by an average of 1.97%, 9.89%, 2.85%, 1.83%, and 1.00% on ODS, OIS, P, F1, and mIoU respectively. Although SimCrack [27] ranked second on most metrics, its actual segmentation performance on fine crack pixels and crack tips was unsatisfactory, as shown in Fig. 6. Other methods such as UCTNet [8] and DTrCNet [7] often failed to distinguish crack regions from background noise or missed crucial fine-grained crack details.
On the more challenging GAPS509 dataset, our method also achieved the highest performance on all metrics except recall. SimCrack obtained a higher recall value of 0.6997, about 2.72% higher than our method, but its precision was significantly lower at 0.6171, leading to a lower F1 score overall. Our method maintained a superior balance between precision and recall, with improvements of 1.31%, 1.25%, 2.90%, 1.11%, and 0.45% on ODS, OIS, P, F1, and mIoU, respectively. Visual results in Fig. 6 further confirm that our method effectively suppresses false detections in irrelevant background areas while preserving accurate segmentation of slender cracks.
Figure 6: Visualization results of CrackSCF and other methods on TUT and other public dataset. Key areas are highlighted with red boxes to indicate easily missed crack pixels and yellow boxes to indicate easily misdetected non-crack areas.
Both DeepCrack and Crack500 datasets are characterized by relatively wide crack regions, testing the model’s capability to capture extensive crack areas. As listed in Table 2, our method achieved the best results across all metrics on the DeepCrack dataset, outperforming the second-best methods by 1.08%, 0.39%, 1.51%, 1.04%, 1.42%, and 0.85% in terms of ODS, OIS, P, R, F1, and mIoU. Although CTCS [9] also performed well, it fell short in preserving crack detail segmentation, as illustrated in Fig. 6. This demonstrates the advantage of our method in accurately segmenting both wide crack regions and their finer structures. On the Crack500 dataset, our method achieved leading performance in five of the six evaluation metrics: ODS, OIS, P, F1, and mIoU, surpassing the second-best methods by 1.11%, 0.07%, 1.69%, 0.47%, and 0.64%. While DTrCNet [7] achieved a higher recall, our method outperformed it in precision by a notable 10.51%, leading to the highest F1 score and indicating a superior balance between detecting true crack pixels and suppressing false positives. Furthermore, as shown in Fig. 6, our method produces fewer erroneous detections in irrelevant background areas and achieves more accurate and refined segmentation of complex crack structures compared to competing approaches.
Table 3: Comparison of the complexity of CrackSCF with other methods
Methods FLOPs ↓ Params ↓ FPS ↑
RIND 695.77G 59.39M 11
UCTNet 171.77G 66.22M 12
SFIAN 52.36G 13.63M 35
CTCS 39.47G 22.88M 28
DTrCNet 123.20G 63.45M 47
Crackmer 14.94G 5.90M 31
SimCrack 70.66G 29.58M 29
Ours 9.26G 4.79M 36
Table 4: Comparison of CrackSCF using LRDS blocks with networks using other convolutions
Method ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
Original Conv 0.8145 0.8193 0.8294 0.8398 0.8346 0.8437 90.06G 17.86M 23
DS Conv 0.8117 0.8166 0.8239 0.8364 0.8301 0.8417 23.91G 8.38M 35
Low-Rank Conv 0.8154 0.8220 0.8220 0.8451 0.8334 0.8441 21.38G 4.63M 30
LRDS Conv 0.8202 0.8266 0.8282 0.8484 0.8382 0.8473 9.26G 4.79M 36
Table 5: Comparison of SCFM performance with other segmentation heads
SegHead ODS OIS P R F1 mIoU FLOPs ↓ Params ↓ FPS ↑
PID 0.7120 0.7198 0.7921 0.6742 0.7284 0.7777 36.69G 5.44M 21
SegFormer 0.7446 0.7634 0.7533 0.7786 0.7658 0.7983 29.72G 6.20M 23
UNet 0.7836 0.7951 0.7829 0.8248 0.8033 0.8222 40.59G 12.34M 20
Hamburger 0.7447 0.7609 0.7075 0.8328 0.7670 0.7969 168.23G 7.45M 14
SCFM 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
Table 6: Comparison of the performance of using our lightweight deformable attention in LDE with other variants
Attention ODS OIS P R F1 mIoU FLOPs Params FPS↑
Common MHA 0.7919 0.8009 0.7999 0.8240 0.8118 0.8279 15.98G 5.80M 16
MHA + LR_Linear 0.7924 0.8003 0.7988 0.8284 0.8134 0.8286 8.87G 4.77M 14
Common MSDA 0.8058 0.8151 0.8052 0.8467 0.8257 0.8373 16.37G 5.82M 28
MSDA + LR_Linear 0.8202 0.8266 0.8289 0.8484 0.8382 0.8473 9.26G 4.79M 36
4.4 Complexity Analysis
To evaluate our LRDS block’s efficacy, we analyzed its computational and parameter efficiency using three metrics: floating point operations (FLOPs), number of parameters (Params), and frames per second (FPS). FLOPs quantify computational load, reflecting demand during inference. Params measure storage requirements, reflecting trainable and stored parameters. FPS gauges real-time processing, indicating frames processed per second under specific hardware conditions.
Table 3 presents our method’s performance alongside other comparisons in terms of FLOPs, Params, and FPS. Our approach achieves notably low values in FLOPs at 9.26G and Params at 4.79M, surpassing the second-best by 38.01% and 18.81%, respectively. This underscores our algorithm’s efficiency, making it suitable for implementation on resource-constrained edge devices. In FPS, our model achieves the second-best performance, processing 36 frames per second, trailing only behind the speed-focused DTrCNet [7]. This highlights our model’s robust real-time capability, crucial for prompt crack image segmentation in practical applications. Overall, our model combines low parameters and computational load with competitive processing speed, ensuring effective deployment on devices like smartphones, drones, and other portable platforms.
4.5 Ablation Studies
We conducted ablation studies on the TUT dataset to assess the effectiveness of our proposed module.
1) Ablation study of different convolution types: As list in Table 4, on six metrics, our method outperforms methods that use depth-separable convolution or simple low-rank approximate convolution. Our method reduces FLOPs by 61.26% and 56.69%, respectively, and Params by 0.16M compared to the simple low-rank approximation convolution, achieving the highest FPS. Compared to original convolution operations, our method reduces FLOPs and Params by 89.70% and 73.19%, respectively, without degrading performance. These results indicate that LRDS effectively reduces computational load and utilizes parameters while maintaining or improving segmentation performance, enhancing processing speed and benefiting lightweight model design.
2) Ablation study of different segmentation heads: To demonstrate the advantages of SCFM over other high-performance segmentation heads, we conducted ablation experiments on different segmentation heads, including PID [28], SegFormer [6], UNet [29], and Hamburger head [30], using the TUT dataset. As list in the Table 5, SCFM achieves the best results in segmentation performance metrics. Particularly in terms of F1 and mIoU, SCFM outperforms the UNet head, which achieved the second-best results, by 4.35% and 3.05%, respectively. This indicates that SCFM effectively enhances the model’s segmentation performance by perceiving and integrating local details and long-range dependencies of crack pixels. Furthermore, SCFM requires low computational resources, with Params and FLOPs being 11.95% and 68.84% lower than those of the second-best results, respectively, and achieves the highest FPS. This demonstrates that our SCFM maintains excellent performance while requiring low computational resources.
3) Ablation study of different attention variants in LDE: To validate that our LR Linear can effectively reduce the computational resources required by the deformable attention (MSDA) in LDE while achieving excellent segmentation performance, we compared lightweight MSDA with common multi-head attention (MHA), MHA with embedded LR Linear, and common MSDA. As shown in the Table 6, our method achieves the best results in segmentation performance metrics. Specifically, compared to the common MSDA, our method improves by 1.51% and 1.19% in F1 and mIoU, respectively. Compared to MHA, MSDA outperforms it in segmentation performance, demonstrating the effectiveness of our use of MSDA. Furthermore, whether for MHA or MSDA, when embedded with LR Linear, the required Params and FLOPs significantly decrease. For MSDA, the Params and FLOPs are reduced by 43.43% and 17.70%, respectively. This demonstrates that our LR Linear can effectively reduce the computational requirements needed by the attention mechanism.
5 Conclusion
This paper proposes a staircase cascaded fusion crack segmentation network (CrackSCF) for pixel-level crack detection. In CrackSCF, we propose the lightweight convolutional block LRDS to construct the MFE, which dramatically reduces the amount of computation and parametric quantities required by the model and is able to efficiently capture the local details of the crack image, and we propose LDE to more efficiently perceive the pixel irregular long-range dependencies in the crack feature maps, in order to more comprehensively fuse the local patterns with the pixel’s long-range dependencies, we also propose SCFM to fully perceive them to generate higher quality segmentation maps. Furthermore, to comprehensively evaluate our model’s robustness and stability across various complex scenarios, we collected the TUT dataset, which includes eight types of crack scenes. Experimental results show that our method performs best on five public datasets and the TUT dataset, particularly excelling in background noise suppression and fine crack segmentation. The computational load and parameter count are reduced to 9.26G and 4.79M, respectively.
Despite our method achieves the low parameter count and computational load, there is still room for improvement in inference speed. In future work, we will further pursue network lightweighting to reduce computational resource requirements and achieve higher inference speeds, aiming to deploy our model on mobile devices for better real-world applications. Additionally, we plan to continue designing an improved SCFM to better fuse and enhance local information and global context for superior segmentation results.